OpenStudy (anonymous):

4. Darnell is selecting names from a hat to determine who will choose the family activities for the weekend. The names in the hat include his mother, father, sister, and him. Calculate the probability of each compound event.

OpenStudy (anonymous):

wait.... will put it on

OpenStudy (anonymous):

Match Term Definition 

OpenStudy (johnweldon1993):

Hmm well 1 by 1 P(Darnell and sister) with replacement there are 4 cards....and when being replaced after each time....the probability stays the same.... so P(darnell) = 1/4 replace P(sister) = 1/4 still so 1/4 * 1/4 = 1/16 P(Mother and Father) without replacement P(mother) = 1/4 not replacing so only 3 cards now P(father) = 1/3 so 1/4 * 1/3 = 1/12 can you continue from here?

OpenStudy (anonymous):

i'll try

OpenStudy (johnweldon1993):

Well do it along with me P(Parent and sister) without replacement P(Parent) how many parent cards are there?

OpenStudy (anonymous):

2

OpenStudy (johnweldon1993):

Right...so P(Parent) = 2/4 now without replacing...there are now only 3 cards to choose from P(sister) how many sister cards are there?

OpenStudy (anonymous):

1

OpenStudy (johnweldon1993):

Correct again....so P(sister) now = 1/3 so to find the compound probability you multiply the 2 2/4 * 1/3 = 2/12...simplified down = 1/6 right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Now....*even though there's only 1 answer left* try that last one P(Parent and Child) with replacement

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when it says with replacement, the child part, what would be that number?

OpenStudy (johnweldon1993):

So after the P(Parent) which is 2/4 right? the card they picked....got put back in...so the total number of cards didn't change...still 4 possible cards.... so P(Child) is.....some number/4 what is some number?

OpenStudy (anonymous):

2/4*2/4=4/16 4/16 divide by 4 = 1/4

OpenStudy (anonymous):

Am I right?

OpenStudy (johnweldon1993):

You are correct! Great job!

OpenStudy (anonymous):

thank you trying to finish the last of math.

OpenStudy (johnweldon1993):

No problem!